OpenStudy (anonymous):

solve this by elimination method x/2+y/5=14/15 and x/3-y/12=2/3 with complete solution please.tnx

OpenStudy (anonymous):

\[\frac{x}{2}+\frac{y}{5}=\frac{14}{15}\] \[\frac{x}{3}-\frac{y}{12}=\frac{2}{3}\]

OpenStudy (anonymous):

working on it now.

OpenStudy (anonymous):

hating fractions, i would multiply everything by 60 before i started messing around \[15x+15y=28\] \[20x-5y=40\] now at least i can see what i am doing

OpenStudy (anonymous):

wait 73 x/2 * 60 = 30x

OpenStudy (anonymous):

looks like multiply second equation by 3 will do it \[30x+15y=56\] \[60x-15y=120\]

OpenStudy (anonymous):

yeah i messed up top one twice actually. it is early. think it s right now yes?

OpenStudy (anonymous):

top one is right.

OpenStudy (anonymous):

\[60\times \frac{14}{15}=14\times 4=56\] i think that line is now correct yes?

OpenStudy (anonymous):

got it from here?

OpenStudy (anonymous):

that is correct., and for bot i just multpied x/3 and 2/3 by 4. so i got 4x - y = 12

OpenStudy (anonymous):

wait i find the LCD THE LCD OF FIRST IS 10 AND THE OTHER 1 IS 12 but i do't know what is the next step

OpenStudy (anonymous):

wait wait, i made an error, its 4x - y = 8

OpenStudy (anonymous):

i had the probably unnecessary large equations \[30x+15y=56\] \[20x-5y=40\]

OpenStudy (anonymous):

yup so i made it 30x +15y = 56 60x -15 = 120

OpenStudy (anonymous):

then say multiply second equation by 3 so you will have 15y and -15y which will add to get 0

OpenStudy (anonymous):

-15y

OpenStudy (anonymous):

@daverer is that step clear? you want to "eliminate" one of the variables

OpenStudy (anonymous):

176/90 = x x = 1.96

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

actually \[x=\frac{176}{90}=\frac{88}{45}=1.9\overline 5\]

OpenStudy (anonymous):

then i sub in for x, and i got y = -0.178

OpenStudy (anonymous):

but rounded you get what you wrote

OpenStudy (anonymous):

this problem is a pain

OpenStudy (anonymous):

and thats rounding its 1.777777

OpenStudy (anonymous):

how do you write it mathematically?

OpenStudy (anonymous):

hold on. may be some mistake here. let me see if i can find it

OpenStudy (anonymous):

... this problem is really a headache ...... this is our new lesson and it's hard.

OpenStudy (anonymous):

just tried machine and got different answer so i will go slow first equation is \[\frac{x}{2}+\frac{y}{5}=\frac{14}{15}\] multiply by 60 to get \[30x+12y=56\] oh heavens i messed this all up. should be \[30x+12y=56\] now what i wrote earlier.

OpenStudy (anonymous):

i wrote it as y/4 that was my fault :(

OpenStudy (anonymous):

ok now we can continue. second equation is \[\frac{x}{3}-\frac{y}{12}=\frac{2}{3}\] multiply by 60 to get \[20x-5y=40\]

OpenStudy (anonymous):

now at this point i would use substitution rather than elimination because second equation is \[4x-y=8\] so you could put \[y=4x-8\] and solve

OpenStudy (anonymous):

were do you multiply it anyway?

OpenStudy (anonymous):

sure lets finish this up. first equation is \[15x+6y=28\] second is \[4x-y=8\]

OpenStudy (anonymous):

multiply second by 6 to get \[15x+6y=28\] \[24x-6y=48\]

OpenStudy (anonymous):

add the two equations to get \[39x=76\]

OpenStudy (anonymous):

and finally \[x=\frac{76}{39}\] a painful but correct answer

OpenStudy (anonymous):

oh.oh it's really hard for me

OpenStudy (anonymous):

yes well this was an annoying one. especially since we got off to a false start. steps were these a) multiply first equation by 15 b) multiply second equation by 12 this clears the fractions c) multiply second equation by 6 so the y's will drop out when you add d) add now you have one equation in one variable. solve e) substitute back to see what the other variable is

OpenStudy (anonymous):

correct answer is \[x=\frac{76}{39},y=-\frac{8}{39}\]

OpenStudy (anonymous):

do you know how to check this? checking..

OpenStudy (anonymous):

yes put the numbers x and y into the other equation as well to see if you get the correct equation.